OpenStudy (anonymous):

Solve. Assume the exercise describes a linear relationship. The average value of a certain type of automobile was $14,040 in 1994 and depreciated to $4140 in 1999. Let y be the average value of the automobile in the year x, where x = 0 represents 1994. Write a linear equation that models the value of the automobile in terms of the year x.

OpenStudy (anonymous):

We need slope and y-intercept. we can use the points (0, 14040) and (5, 4140) How do we find slope?

OpenStudy (anonymous):

a. y = -1980x - 5760 b. y = -1980x + 4140 c. y = - 1/1980x - 4140 d. y = -1980x + 14,040

OpenStudy (anonymous):

Are you okay with (0, 14040) and (5, 4140)? How do we find slope?